# AUTALIC: A Dataset for Anti-AUTistic Ableist Language In Context

### v2

@@ Abstract
As our understanding of autism and ableism continues to increase, so does our understanding of ableist language towards autistic people. Such language poses a significant challenge in NLP research due to its subtle and context-dependent nature. Yet, detecting anti-autistic ableist language remains underexplored, with existing NLP tools often failing to capture its nuanced expressions. We present AUTALIC, the first benchmark dataset dedicated to the detection of anti-autistic ableist language in context, addressing a significant gap in the field. The dataset comprises 2,400 autism-related sentences collected from Reddit, accompanied by surrounding context, and is annotated by trained experts with backgrounds in neurodiversity. Our comprehensive evaluation reveals that current language models, including state-of-the-art LLMs, struggle to reliably identify anti-autistic ableism and align with human judgments, underscoring their limitations in this domain. We publicly release AUTALIC along with the individual annotations which serve as a valuable resource to researchers working on ableism, neurodiversity, and also studying disagreements in annotation tasks. This dataset serves as a crucial step towards developing more inclusive and context-aware NLP systems that better reflect diverse perspectives.
@@
Trigger warning: this paper contains examples of anti-autistic ableist language including explicit slurs and references to violence.
1 Introduction
@@
There are several critical frameworks used to define and understand disability, including autism Lawson and Beckett (2021). While these frameworks are ever-evolving, the medical model framework which defines disability as a “disease” is one of the most commonly applied in computer science research when discussing disability in general and autism specifically Rizvi et al. (2024); Spiel et al. (2019a); Sideraki and Drigas (2021); Anagnostopoulou et al. (2020); Parsons et al. (2020); Williams et al. (2023); Sum et al. (2022). This framework defines autism as a deficit of skills such as the Theory of Mind, and its applications in technology research largely focus on providing diagnosis and treatment to autistic people Baron-Cohen (1997); Begum et al. (2016); Rizvi et al. (2024); Spiel et al. (2019b). However, this framework has been criticized for creating a power imbalance between medical professionals and the autistic community as it promotes the belief that a formal diagnosis is necessary to validate the autistic identity Bennett and Keyes (2020). This belief also posits neurotypical behaviors as the “norm” and autism as a “deficit” of these norms, thereby promoting neuronormativity instead of neurodiversity, which views all neurotypes as valid forms of human diversity Bottema-Beutel et al. (2021); Walker (2014).
In contrast, prior work has shown viewing autism as an identity may increase the psychological well-being of autistic individuals and lower their social anxiety Cooper et al. (2023). Although the language used to describe autism varies, prior studies with autistic American adults found 87% prefer identity-first language over person-first language Taboas et al. (2023). Person-First Language (PFL) centers the person (i.e. person with autism), while Identity-First Language (IFL) centers the identity (i.e. autistic person) Taboas et al. (2023).
Prior work examining ableist speech directed at content creators identified 11 types of “ableist hate and harassment" that they categorized as slurs and derogatory language, violent and eugenics-related speech, questioning ability and denying access, mocking and invalidation of disability identity, and objectifying the disabled body Heung et al. (2024). Additionally, researchers have identified emerging terms such as “weaponized autism" that are gaining popularity on alt-right social media platforms, further reflecting the ever-evolving nature of negative societal attitudes towards autism Welch et al. (2023).
@@
Anti-autistic ableist language can be diverse in scope. It may include perpetuating stereotypes, using offensive language and slurs, or centering non-autistic people over the perspectives of autistic people Bottema-Beutel et al. (2021); Rizvi et al. (2024); Darazsdi and Bialka (2023). While abusive language detection systems can help identify such speech, they are known to demonstrate bias Manerba and Tonelli (2021); Venkit et al. (2022), with even LLMs perpetuating ableist biases Gadiraju et al. (2023a). Additionally, anti-autistic ableist speech remains understudied, which is concerning given that classifiers trained on multiple hate speech datasets have shown a failure to generalize to target groups outside of the training corpus Yoder et al. (2022).
Additionally, no datasets specifically focus on anti-autistic speech classification, and only 3 of the 23 datasets for bias evaluation in LLMs focus on disability Gallegos et al. (2024). Prior work has also found that LLMs lack an acknowledgment of context, which leads to higher rates of false positives when classifying ableist speech Phutane et al. (2024). These limitations are also found in toxicity classifiers, which excel primarily at identifying explicit ableist speech but may otherwise perpetuate harmful social biases leading to content suppression Phutane et al. (2024).
Our work addresses these gaps in existing research by examining the performance of our dataset in the following domains for the task of anti-autistic hate speech classification:
@@
    The impact our dataset has on fine-tuning classifiers

### v3

@@ Abstract
As our understanding of autism and ableism continues to increase, so does our understanding of ableist language towards autistic people. Such language poses a significant challenge in NLP research due to its subtle and context-dependent nature. Yet, detecting anti-autistic ableist language remains underexplored, with existing NLP tools often failing to capture its nuanced expressions. We present Autalic, the first dataset dedicated to the detection of anti-autistic ableist language in context, addressing a significant gap in the field. Autalic comprises 2,400 autism-related sentences collected from Reddit, accompanied by surrounding context, and annotated by trained experts with backgrounds in neurodiversity. Our comprehensive evaluation reveals that current language models, including state-of-the-art LLMs, struggle both to reliably identify anti-autistic ableism and to align with human judgments, underscoring their limitations in this domain. We publicly release Autalic along with the individual annotations. This dataset serves as a crucial step towards developing more inclusive and context-aware NLP systems that better reflect diverse perspectives.
@@
Trigger warning: this paper contains ableist language including explicit slurs and references to violence.
1 Introduction
@@
There are several critical frameworks used to define autism Lawson and Beckett (2021), including the medical model, which defines disability as a “disease” and is one of the most widely used in computer science research focusing on autism Rizvi et al. (2024); Spiel et al. (2019a); Sideraki and Drigas (2021); Anagnostopoulou et al. (2020); Parsons et al. (2020); Williams et al. (2023); Sum et al. (2022).
@@
Since this framework defines autism as a deficit of skills, its applications in technology research largely focus on providing diagnosis and treatment to autistic people Baron-Cohen (1997); Begum et al. (2016); Rizvi et al. (2024); Spiel et al. (2019b). This belief also posits neurotypical behaviors as the “norm” and autism as a “deficit” of these norms, thereby promoting neuronormativity instead of neurodiversity, which views all neurotypes as valid forms of human diversity Bottema-Beutel et al. (2021a); Walker (2014).
To improve the alignment of AI research with neurodiversity, we present Autalic, a dataset of 2,400 autism-related sentences that we collect from various communities on Reddit, along with the original context of 2,014 immediately preceding and 2,400 following sentences. We aim to fill a critical gap in current NLP research, which has largely overlooked the nuanced and context-dependent nature of ableist speech targeting autistic individuals. Our dataset not only captures key contextual elements but also incorporates a comprehensive annotation process led by trained annotators with an understanding of the autistic community, ensuring higher reliability and relevance. Our final dataset contains all of the labels to capture the nuances in human perspectives, and allows Autalic to serve as a resource for researchers studying anti-autistic ableist speech, neurodiversity, or disagreements in general.
Through a series of experiments with classical models and 4 LLMs, we find empirical evidence highlighting the difficulty of this task, and that LLMs are not reliable agents for such annotations. Our evaluations indicate that reasoning LLMs have the most consistent scores regardless of the language used in the prompt, thereby indicating a more thorough understanding of the different ways anti-autistic speech may be identified or may manifest in text. We find that in-context learning examples provide mixed results in helping improve the task comprehension among LLMs.
2 Related Work
@@
Anti-autistic ableist language can be diverse in scope. It may include perpetuating stereotypes, using offensive language and slurs, or centering non-autistic people over the perspectives of autistic people Bottema-Beutel et al. (2021a); Rizvi et al. (2024); Darazsdi and Bialka (2023). While abusive language detection systems can help identify such speech, they are known to demonstrate bias Manerba and Tonelli (2021); Venkit et al. (2022), with even LLMs perpetuating ableist biases Gadiraju et al. (2023a). Additionally, anti-autistic ableist speech remains understudied, which is concerning given that classifiers trained on multiple hate speech datasets have shown a failure to generalize to target groups outside of the training corpus Yoder et al. (2022).
Although the language used to describe autism varies, prior studies with autistic American adults found 87% prefer identity-first language over person-first language Taboas et al. (2023). Person-First Language (PFL) centers the person (e.g. “person with autism”), while Identity-First Language (IFL) centers the identity (e.g. “autistic person”) Taboas et al. (2023). Supporting this finding, other researchers have found that viewing autism as an identity may increase the psychological well-being of autistic individuals and lower their social anxiety Cooper et al. (2023).
Ableist language online varies, may manifest in different ways and is ever-evolving  Heung et al. (2024); Welch et al. (2023). However, toxic language datasets focusing on hate speech and abusive language have often addressed disability in general terms but have not explicitly focused on autism ElSherief et al. (2018); Ousidhoum et al. (2019). To our knowledge, there are no previous datasets specifically focused on anti-autistic speech classification, and only 3 of the 23 datasets for bias evaluation in LLMs focus on disability Gallegos et al. (2024). LLMs may be limited in that they lack an acknowledgment of context, which leads to higher rates of false positives when classifying ableist speech Phutane et al. (2024). These limitations are also found in toxicity classifiers, which excel primarily at identifying explicit ableist speech but may otherwise perpetuate harmful social biases leading to content suppression Phutane et al. (2024). Toxic language detection models, including LLM-based models, have been found to exhibit strong negative biases toward disabilities by classifying any disability-related text as toxic Narayanan Venkit et al. (2023). Further, LLMs have been observed to perpetuate implicitly ableist stereotypes Gadiraju et al. (2023b) and bias Gama (2024); Venkit et al. (2022). This, unfortunately, can sometimes be due to a research design that overlooks intra-community and disabled people’s perspectives Mondal et al. (2022), as well as autistic people’s views, which may lead to harmful stereotypes Rizvi et al. (2024); Spiel et al. (2019b). We make a step towards addressing these issues by building a dataset that focuses on ableist speech and autism by including autistic people’s perspectives during the annotation processing as recommended by Davani et al. (2023). Autalic contains all its labels and will also be useful for researchers interested in leveraging disagreements for difficult classification tasks Leonardelli et al. (2021); Pavlick and Kwiatkowski (2019).
3 Autalic
To build Autalic, we collected relevant sentences containing autism-related keywords from Reddit using the methods described in Section 3.1. The collected sentences were labeled by trained annotators, as discussed in Section 3.2.
3.1 Data Collection
We identify a methodology for curating sentences related to autism similar to prior datasets by collecting English-language sentences from Reddit  Overbay et al. (2023); Wadhwa et al. (2023); Njoo et al. (2023); Antoniak et al. (2024); Sabri et al. (2024); Allein and Moens (2024); Buz et al. (2024). The limitations of this method are detailed in Section  6.
3.1.1 Data Collection Criteria
We select Reddit as our source based on its popularity, focus on text-based content, and fewer API restrictions than X at the time of our data collection in January 2024. We search for keywords using the default search settings, which filters posts based on relevancy by prioritizing rare words in the search query, the age of the post, and the amount of likes and comments it has.  [^1] https://support.reddithelp.com/hc/en-us/articles/19695706914196-What-filters-and-sorts-are-available The search terms include “autis*”, “ASD”, “aspergers”, and “disabilit*”; the full list is available in the Appendix, Table 4.
We use the identified search terms to collect the target sentence instance containing our keywords, to be labeled by the annotators, and the sentences preceding or following target instances to provide additional context. We collect 2,400 target sentences, with 2,014 preceding and 2,400 following them. Finally, we split our dataset into three parts by randomly selecting and assigning 800 unique target sentences to create three segments that were each annotated by a group of three annotators.
The average number of likes on each post included in the Autalic dataset is 1,611.59. Table 1 details the subreddits from which the most significant number of sentences were extracted from. With the exception of r/AmITheAsshole, all of the other subreddits are autism-related.
SubReddit Sentence Count
r/Aspergers 116
r/Autism 88
r/AmITheAsshole 39
r/AutisminWomen 37
r/AuDHDWomen 24
3.1.2 Data Curation
As some of the identified keywords may appear in other contexts, we perform an exact word search for the acronyms to ensure unrelated words that might contain our acronyms are excluded from the search as they go beyond the scope of our dataset. For example, we searched for “applied behavioral analysis” and a case-sensitive search for “ABA”, which is a form of therapy intended to minimize autistic behaviors such as stimming (which is often used for self-soothing) Sandoval-Norton et al. (2019). Similarly, we exclude any posts that are not written in English using the Python package langdetect and posts that contain images, videos, or links.  [^2] https://pypi.org/project/langdetect/
3.1.3 Final Dataset
Our final dataset includes 2,400 sentences from 192 different subreddits. To protect our annotators’ privacy, we have anonymized individual label selections.
3.2 Data Annotation
3.2.1 Annotator Selection
We recruit nine upper level undergraduate researchers as volunteer annotators and randomly assign them to annotate different segments of the dataset. We ensure that their involvement in our annotation process is voluntary, informed, and mutually beneficial. In particular, some participants choose to volunteer because they care deeply about the subject matter and wish to contribute to improving AI for autistic people. We select participants for whom this collaboration would provide relevant and valuable professional experience, grant them opportunities to engage in other aspects of the research, and provide mentorship and authorship recognition in accordance with the ACL guidelines.
We prioritize the well-being and autonomy of our annotators by providing full disclosure of the research process and subject material prior to their participation. We supply relevant trigger warnings, discuss the nature of the content in detail during an orientation session, and allow annotators to make an informed decision about whether they wish to proceed. We make them aware that they are free to withdraw from the study at any point.
Our annotators are US-based, culturally diverse, and include people who grew up outside the United States. They are all fluent in English. Four of our annotators are gender minorities, and at least three self-identify as neurodivergent. Although we ensure the annotators were from diverse backgrounds during our recruitment process, due to the collaborative nature of our annotation process, we do not share the individual details of their identities. We also note that any personally identifiable information was destroyed upon the conclusion of our analysis and not shared outside of our research team.
3.2.2 Annotator Training
We provide a virtual orientation to all annotators explaining the history of anti-autistic ableism, examples of contemporary anti-autistic discrimination, and a brief overview of the annotation task.
To explain the annotation task, we provide examples of sentences similar to what they may encounter while annotating. For example, we discuss how the inclusion of “at least” alters the connotations of the following sentence:
At least I am not autistic.
We also introduce our glossary to the annotators as a dynamic resource that can be altered as needed. This glossary contains words that may appear in autism discourse online that may not be commonly known to others. These include medical acronyms, slang, and references to organizations and resources commonly affiliated with the autistic community (such as Autism Speaks). An excerpt of our glossary is available in our Appendix. We conclude our orientation by providing a brief tutorial video demonstrating how to run the script that will guide each annotator through the annotation task.
3.2.3 Data Labeling
After completing the training, we assign each of the three segments of the dataset to three randomly selected annotators. Each annotator is assigned 800 unique sentences, with a goal of completing 200 annotations each week over four weeks. Annotators select from three possible labels for each sentence: “Ableist,” “Not Ableist,” or “Needs More Context.”
Ableist (1): We ask our annotators to select this label if a sentence contains ableist sentiments as defined by the Center for Disability Rights: “Ableism is a set of beliefs or practices that devalue and discriminate against people with physical, intellectual, or psychiatric disabilities and often rests on the assumption that disabled people need to be ‘fixed’ in one form or the other.” [^3] https://cdrnys.org/blog/uncategorized/ableism/
Not Ableist (0): Annotators select this label for sentences that describe positive or neutral behaviors and attitudes regarding autism, or posts written by an autistic person reaching out for help and support. This includes individuals using medical terminology in a personal context (e.g., “I need therapy”), intra-community discussions, and general discussions of medical processes (unrelated to neurodivergence). Some examples of statements labeled as not ableist are: “I am autistic”, “As an autistic person, I think…”
Needs More Context (-1): This label is used for sentences an annotator is unable to definitively categorize as ableist or not ableist even with the contextual sentences provided. This category includes text that is entirely unrelated to disabilities or remains ambiguous without additional context.
The number of times our annotators assign each label is detailed in Table 2. To calculate our agreement scores, we consolidated labels -1 and 0 together based on feedback from our annotators that unrelated sentences needing more context could be classified as not anti-autistic in a purely granular classification.
Label Definition Count
-1 unrelated to autism or needs more context 595
0 not ableist 5,582
1 ableist 1,023
Table 2: An explanation of the labels used in our classification task and the resulting counts of each label from all 9 annotators combined.
While we use the majority label as the ground truth in our analysis, in our public dataset, we will be releasing the individual labels from each annotator due to a growing interest in embracing disagreements for such classification tasks in NLP Leonardelli et al. (2021); Kralj Novak et al. (2022); Pavlick and Kwiatkowski (2019); Plank (2022); Plank et al. (2014).
3.2.4 Providing Context
3.2.5 Disagreements
The average Fleiss’s Kappa scores are 0.25. This score underlines the difficulty of our classification task, which is apparent from the findings of prior works Ousidhoum et al. (2019), including a quantitative assessment of tag confusions that found the majority of disagreements are due to linguistically debatable cases rather than errors in annotation  Plank et al. (2014). Examples of such cases are provided in our Appendix.
We analyze the sentences with the highest levels of disagreement in our dataset. In 100 of these posts, we observe:
@@ 3.2.5 Disagreements
    a tendency to use the medical model terminology or stereotypes (n=48)